The joint density function of two random variables X and Y is given by


1
Expert's answer
2021-09-07T16:39:03-0400

Let use the following problem.



The marginal pdf and expectations are as follows:

"p(x) = \int_y p(x,y) \\

= \int^5_1 \frac{xy}{96} dy \\

= \frac{x}{96} \int^5_1 ydy \\

= \frac{x}{96} [\frac{y^2}{2}]^5_1 \\

= \frac{x}{96}[\frac{25}{2}-\frac{1}{2}] \\

= \frac{x}{8} \; 0<x<4"

The expectation of X is as follows:

"E(X) = \int^4_0x \times \frac{x}{8} dx \\

= \frac{1}{8} \int^4_0 x^2 dx \\

=\frac{1}{8}[\frac{x^3}{3}]^4_0 \\

=\frac{1}{8} \times \frac{4^3}{3} \\

= \frac{8}{3} \\

= 2.6667 \\

p(y) = \int_4 p(x,y) \\

= \int^4_0 \frac{xy}{96} dx \\

= \frac{y}{96} \int^4_0 x dx \\

= \frac{y}{96} [\frac{x^2}{2}]^4_0 \\

= \frac{y}{96} [\frac{16}{2}] \\

= \frac{y}{12} \; 1<y<5 \\

E(y) = \int^5_1 y \times \frac{y}{12}dy \\

= \frac{1}{12} \int^5_1 y^2 dx \\

= \frac{1}{12}[\frac{y^3}{3}]^5_1 \\

= \frac{1}{12} \times \frac{5^3 -1}{3} \\

= \frac{31}{9} \\

= 3.4444 \\

p(x) \times p(y) = \frac{x}{8} \times \frac{y}{12} \\

=\frac{xy}{96} \\

=p(x,y)"

Thus, X and Y are independent.

"E(XY)=E(x) \times E(y) \\

= \frac{8}{3} \times \frac{31}{9} \\

= \frac{248}{27} \\

= 9.1852 \\

E(2X+3Y) = 2E(x) +3E(y) \\

= 2 \times \frac{8}{3} + 3 \times \frac{31}{9} \\

= \frac{16}{3} + \frac{31}{3} \\

= \frac{47}{3} \\

= 15.6667"